# Collapsing Flat ${\rm{SU}}(2)$-Bundles to Spherical $3$-Manifolds

### 17 Mar 2026 v1

@@ Abstract
We present a geometric mechanism for the emergence of spherical $3$-manifolds from the superspace of Riemannian metrics associated with flat ${\rm{SU}}(2)$-bundles over closed orientable hyperbolic surfaces. Our main result shows that any spherical 3-manifold $(S,g_{S})$ can be realized as a boundary point in the Gromov-Hausdorff closure of a superspace $\mathcal{S}(P)$, where $P$ is a flat ${\rm{SU}}(2)$-bundle over a closed orientable hyperbolic surface $(\Sigma,h_{\Sigma})$. We show that the convergence of the sequence of metric spaces towards the spherical limit is controlled by the order of the fundamental group of $S$ and the metric invariant of the hyperbolic base provided by the ratio between its area and its systole. In this framework, the problem of obtaining the sharpest upper bound error reduces to the classical problem of maximizing the systole function over the moduli space of hyperbolic Riemann surfaces. As a byproduct, we observe that certain arithmetic surfaces provide the best possible error estimates within this family. To illustrate these results, we show that, according to our mechanism, the Bolza surface yields the optimal error bound for the convergence toward the Poincaré homology sphere.
@@ 1 Introduction
In Riemannian geometry, the significance of these groups is realized through the spherical $3$-manifolds. By the Killing-Hopf theorem [WOL11, §2.4], a complete, simply connected Riemannian manifold of constant positive curvature +1+1 is isometric to the $3$-sphere $S^{3}\cong{\rm{SU}}(2)$. Consequently, any closed orientable spherical $3$-manifold is isometric to a quotient ${\rm{SU}}(2)/\Gamma_{ADE}$, where $\Gamma_{ADE}\subset{\rm{SU}}(2)$ is a finite subgroup and therefore one of the binary polyhedral groups from the ADE list [SCO83], [MIL75]. These manifolds form a crucial class within Thurston’s geometrization program [THU22b], [THU22a]. The geometrization conjecture, proven by Perelman [PER02, PER03b, PER03a], asserts that every closed $3$-manifold can be decomposed into pieces, each admitting one of eight geometric structures. The spherical manifolds are precisely those that admit a metric of constant positive curvature, serving as the most symmetric models in this classification.
@@
we observe a remarkable collapsing phenomenon. As $n\to\infty$, the diameter of the hyperbolic base shrinks to zero, while the ${\rm{SU}}(2)$ fibers remain unchanged. The Gromov-Hausdorff limit [BBI01] of this family is precisely the spherical $3$-manifold ${\rm{SU}}(2)/\Gamma_{ADE}$ associated with the holonomy group of the flat connection.
Thus, flat ${\rm{SU}}(2)$-bundles over hyperbolic surfaces serve as geometric interpolators between two geometries of opposite curvature signs: the negative curvature of the hyperbolic base (K=1K=-1) and the positive curvature of the spherical fiber ($K=+1$). The controlled collapse erases the hyperbolic geometry, revealing the underlying spherical geometry of the quotient. Our main theorem shows that every orientable spherical $3$-manifold arises as a boundary point in the Gromov–Hausdorff closure of the superspace ([FIS70], [EDW75]) of Riemannian metrics associated with such a flat bundle. Moreover, the rate of this convergence is quantitatively controlled by the order of the fundamental group of the spherical manifold and the metric invariant provided by the ratio between the area and the systole (e.g. [GRO83], [KAT07]) of the hyperbolic base. This establishes a direct geometric link between the topological realization of Thurston’s spherical geometries and the systolic optimization of the moduli space of Riemann surfaces, offering a dynamical perspective on the emergence of ADE-classified 3-manifolds.
@@
Let $(S,g_{S})$ be an orientable spherical $3$-manifold. Then, there exist a closed orientable hyperbolic surface $\Sigma\in\mathcal{M}_{g}$ and $P,\theta]\in\mathcal{M}_{{\text{flat}}}(\Sigma,{\rm{SU}}(2))$, such that
@@
such that $p$ is any odd prime number and $\nu$ depends on a division quaternion algebra $A$, satisfying
@@
where $C$ is a positive constant depending only on $A$. Combining the above result with Theorem A, we have the following corollary.
@@
Let $(S,g_{S})$ be an orientable spherical $3$-manifold. Then, there exist a closed orientable hyperbolic surface $\Sigma$ of genus $g(\Sigma)>1$ and $P,\theta]\in\mathcal{M}_{{\text{flat}}}(\Sigma,{\rm{SU}}(2))$, such that
@@ 3 Flat Principal Bundles
A principal connection on a principal $G$-bundle $G\hookrightarrow P\to M$ is a $\mathfrak{g}$-valued $1$-form $\theta\colon TP\to\mathfrak{g}$ on $P$, satisfying the following
@@
Let $P$ be a principal $G$-bundle over $M$, and let $\theta$ be a principal connection on $P$. We define the curvature $2$-form $D\theta\in\Omega^{2}(P,\mathfrak{g})$ of the connection $\theta$ as
@@
such that $Z:=RST$ is a central element of order $2$, see for instance [COX40], [CM13]. The groups as above are finite if and only if $(a,b,c)$ satisfies
@@ 4 Spherical 3-manifolds and Flat ${\rm{SU}}(2)$-bundles
Under the identification $S^{3}\cong{\rm{SU}}(2)$, from the Killing-Hopf theorem and Synge’s theorem, it follows that a spherical $3$-manifold is isometric to a homogeneous ${\rm{SU}}(2)$-space, namely, we have an isometry
@@ 5 Gromov-Hausdorff Space
Given a metric space $(X,d_{X})$, for every pair of subsets $A,B\subset X$, we have the Hausdorff distance between $A$ and $B$ given by
@@ 7 Proof of Main Result
Given a spherical $3$-manifold $(S,g_{S})$, we have an isometry
@@ 8 From Bolza Surface to Poincaré Homology Sphere
The Bolza surface $\Sigma$ is a Riemann surface of genus $2$ with a holomorphic automorphism group of order 48, the highest for this genus [BOL87], [JEN84], [SCH93]. The surface $\Sigma$ can be viewed as the smooth completion of its affine form

### 18 Mar 2026 v2

# Collapsing Flat ${\rm{SU}}(2)$-Bundles to Spherical 33-Manifolds

@@ Abstract
We present a geometric mechanism for the emergence of spherical 33-manifolds from the superspace of Riemannian metrics associated with flat ${\rm{SU}}(2)$-bundles over closed orientable hyperbolic surfaces. Our main result shows that any spherical 3-manifold $(S,g_{S})$ can be realized as a boundary point in the Gromov-Hausdorff closure of a superspace $\mathcal{S}(P)$, where $P$ is a flat ${\rm{SU}}(2)$-bundle over a closed orientable hyperbolic surface $(\Sigma,h_{\Sigma})$. We show that the convergence of the sequence of metric spaces towards the spherical limit is controlled by the order of the fundamental group of $S$ and the metric invariant of the hyperbolic base provided by the ratio between its area and its systole. In this framework, the problem of obtaining the sharpest upper bound error reduces to the classical problem of maximizing the systole function over the moduli space of hyperbolic Riemann surfaces. As a byproduct, we observe that certain arithmetic surfaces provide the best possible error estimates within this family. To illustrate these results, we show that, according to our mechanism, the Bolza surface yields the optimal error bound for the convergence toward the Poincaré homology sphere.
@@ 1 Introduction
In Riemannian geometry, the significance of these groups is realized through the spherical 33-manifolds. By the Killing-Hopf theorem [WOL11, §2.4], a complete, simply connected Riemannian manifold of constant positive curvature +1+1 is isometric to the 33-sphere $S^{3}\cong{\rm{SU}}(2)$. Consequently, any closed orientable spherical 33-manifold is isometric to a quotient ${\rm{SU}}(2)/\Gamma_{ADE}$, where $\Gamma_{ADE}\subset{\rm{SU}}(2)$ is a finite subgroup and therefore one of the binary polyhedral groups from the ADE list [SCO83], [MIL75]. These manifolds form a crucial class within Thurston’s geometrization program [THU22b], [THU22a]. The geometrization conjecture, proven by Perelman [PER02, PER03b, PER03a], asserts that every closed 33-manifold can be decomposed into pieces, each admitting one of eight geometric structures. The spherical manifolds are precisely those that admit a metric of constant positive curvature, serving as the most symmetric models in this classification.
@@
we observe a remarkable collapsing phenomenon. As $n\to\infty$, the diameter of the hyperbolic base shrinks to zero, while the ${\rm{SU}}(2)$ fibers remain unchanged. The Gromov-Hausdorff limit [BBI01] of this family is precisely the spherical 33-manifold ${\rm{SU}}(2)/\Gamma_{ADE}$ associated with the holonomy group of the flat connection.
Thus, flat ${\rm{SU}}(2)$-bundles over hyperbolic surfaces serve as geometric interpolators between two geometries of opposite curvature signs: the negative curvature of the hyperbolic base (K=1K=-1) and the positive curvature of the spherical fiber ($K=+1$). The controlled collapse erases the hyperbolic geometry, revealing the underlying spherical geometry of the quotient. Our main theorem shows that every orientable spherical 33-manifold arises as a boundary point in the Gromov–Hausdorff closure of the superspace ([FIS70], [EDW75]) of Riemannian metrics associated with such a flat bundle. Moreover, the rate of this convergence is quantitatively controlled by the order of the fundamental group of the spherical manifold and the metric invariant provided by the ratio between the area and the systole (e.g. [GRO83], [KAT07]) of the hyperbolic base. This establishes a direct geometric link between the topological realization of Thurston’s spherical geometries and the systolic optimization of the moduli space of Riemann surfaces, offering a dynamical perspective on the emergence of ADE-classified 3-manifolds.
@@
Let $(S,g_{S})$ be an orientable spherical 33-manifold. Then, there exist a closed orientable hyperbolic surface $\Sigma\in\mathcal{M}_{g}$ and $P,\theta]\in\mathcal{M}_{{\text{flat}}}(\Sigma,{\rm{SU}}(2))$, such that
@@
such that $p$ is any odd prime number and $\nu$ depends on a division quaternion algebra AA, satisfying
@@
where $C$ is a positive constant depending only on AA. Combining the above result with Theorem A, we have the following corollary.
@@
Let $(S,g_{S})$ be an orientable spherical 33-manifold. Then, there exist a closed orientable hyperbolic surface $\Sigma$ of genus $g(\Sigma)>1$ and $P,\theta]\in\mathcal{M}_{{\text{flat}}}(\Sigma,{\rm{SU}}(2))$, such that
@@ 3 Flat Principal Bundles
A principal connection on a principal $G$-bundle $G\hookrightarrow P\to M$ is a $\mathfrak{g}$-valued 11-form $\theta\colon TP\to\mathfrak{g}$ on $P$, satisfying the following
@@
Let $P$ be a principal $G$-bundle over $M$, and let $\theta$ be a principal connection on $P$. We define the curvature 22-form $D\theta\in\Omega^{2}(P,\mathfrak{g})$ of the connection $\theta$ as
@@
such that $Z:=RST$ is a central element of order 22, see for instance [COX40], [CM13]. The groups as above are finite if and only if $(a,b,c)$ satisfies
@@ 4 Spherical 3-manifolds and Flat ${\rm{SU}}(2)$-bundles
Under the identification $S^{3}\cong{\rm{SU}}(2)$, from the Killing-Hopf theorem and Synge’s theorem, it follows that a spherical 33-manifold is isometric to a homogeneous ${\rm{SU}}(2)$-space, namely, we have an isometry
@@ 5 Gromov-Hausdorff Space
Given a metric space $(X,d_{X})$, for every pair of subsets $A,B\subset X$, we have the Hausdorff distance between AA and $B$ given by
@@ 7 Proof of Main Result
Given a spherical 33-manifold $(S,g_{S})$, we have an isometry
@@ 8 From Bolza Surface to Poincaré Homology Sphere
The Bolza surface $\Sigma$ is a Riemann surface of genus 22 with a holomorphic automorphism group of order 48, the highest for this genus [BOL87], [JEN84], [SCH93]. The surface $\Sigma$ can be viewed as the smooth completion of its affine form